OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 10x2 + 24x

OpenStudy (anonymous):

\[ f(x) = x^3 + 10x^2 + 24x = x(x^2+10x+24) \]Clearly one of the zeros is \(x=0\). Use the quadratic formula to find the rest.

OpenStudy (anonymous):

so is it 0,4 & 6

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

On second.

OpenStudy (anonymous):

you got the sign wrong on your zeros.

OpenStudy (anonymous):

Clearly the should all be negative.

OpenStudy (anonymous):

oh alright thnx